OpenStudy (anonymous):

Find the distance between the two points: (7,12), (1,4) i got 10 @Leader @sangya21

OpenStudy (leader):

yes that is correct :)

OpenStudy (leader):

Do you want to see the work?

OpenStudy (anonymous):

ok yes please

OpenStudy (leader):

Because after you plug in everything you get \(\begin{aligned} d(A,B) & = \sqrt{\left( 1 - 7 \right)^2 + \left( 4 - 12 \right)^2} \\ d(A,B) & = \sqrt{ 36 + 64} \\ d(A,B) & = 10 \end{aligned}\)

OpenStudy (anonymous):

oh i get it know thanks

OpenStudy (anonymous):

what about this one Decide whether the points are vertices of a right triangle. (0,0), (0,3), (4,0) a. points are vertices of a right triangle b. points are not vertices of a right triangle i choose a

OpenStudy (leader):



OpenStudy (anonymous):

Check distance between three vertices and see if these distances fit in Pythagoras theorem?

OpenStudy (leader):